the plane PQ X PR 1. Find unit vector the perpendicular to P(1,1,1), Q(2,1,3), R(2, 2, 1).
Answer #1

P1111) & (2, 113) R(2, 2) Pa = (2-1, 14, 3-1) = (0,2) PR = (2-1, 2-1, 1-1)=(1,0) PQ XPR = li (0-2) -(0-2)+K(1-0) -2i+2+kunit perpendicular rector PQ XPRO (POXPRI -zitzitk Haitzitkl -2i+2;+k 14+4+1 2it24 3